A woman sells 7/10 of her stock of cutlery at a profit of 30%. The remaining stock is sold at a loss of 8%. Find the percentage profit on the sale of the whole stock.
Answer by ikleyn(52848)

.
.
A woman sells 7/10 of her stock of cutlery at a profit of 30%.
The remaining stock is sold at a loss of 8%.
Find the percentage profit on the sale of the whole stock.
~~~~~~~~~~~~~~~~

Let x is her bying cost.

For 7/10 of her stock the woman got  0.7*(1+0.3)x = 0.7*1.3x = 0.91x.

For remaining 3/10 of her stock the woman got  0.3*(1-0.08)x = 0.3*0.92x = 0.276x.

In total, the woman got  0.91x + 0.276x = 1.186x.


Her percentage profit is 18.6% of the whole stock.    ANSWER

Solved.